the sample size for a simple random sample from a population are given below x=8,n.32. Ho: p-o4. Ha : p<0.4, α:010 a. Determine the sample proportion. b. Decide whether using the one-proportion z-test is appropriate. c. If appropriate, use the one-proportion z-test to perform the speified hypothesis test. Click here to view a table of areas under the standard normal curve for negative values of z a. The sample proportion is (Type an integer or a decimal. Do not round.) b. Is it appropriate to use the one-proportion z-test? O No c. Determine the test statistic, if appropriate. Select the correct choice below OA (Round to two decimal places as needed.) and, if necessary, fill in the answer box to complete your answer O B. The one-proportion z-test is not appropriate identifly the oritical value(s) if appropriate. Select the correct choice below and, if necessary, ill in the answer box to complete your answer. 2q/2(Round to two decimal places as needed)
The number of successes and the sample size for a simple random sample from a population are given below x-8, n=32, Ho: p-o4. Ha, p < 0.4, α-o.10 a. Determine the sample proportion b. Decide whether using the one-proportion z-test is appropriate. c. If appropriate, use the one-proportion z-test to perform the specified hypothesis test. Cick here to viewa labe.of areas.une the standand noma curve for tivevae of entfly the critical value(s), i appropriate. Select the correct choice below and, f necessary,ill the answer box to complete your answer OA. ±ZW2=堌(Round to two decimal places as needed ) OB. (Round to two decimal places as needed.) OC.(Round to two decimal places as needed) 0 D. The one-proportion z-test is not appropriate. What is the conclusion of the hypothesis test? 0 A, O B. Reject Ho: the data do not provide sufficient evidence to conolude that p<0.4. Reject Ho; the data do provide sufficient evidence to conclude that p <04. C. Do not reject Ho: the data do not provide sufficient evidence to conclude that p<0.4 OD. Donot reject Ho-the data do provide sumcient evidence to conclude that p < 0 4 O E. The one-proportion z-test is not appropriate Click to select your answer(s).
Answer #1

SolutionA

sample proportion=x/n=8/32=0.25

Solutionb:

its a one proportion z test

YES

Solutionc:

ztest statistic

z=p^-p/sqrt(p*(1-p)/n

Z=0.25-0.4/sqrt(0.4*(1-0.4)/32)=-1.73

z=-1.73

ANSWER:

optionA

z=-1.73

Solutiond:

z alpha/2 for 99%=-2.576

answer:

option B

-z alpha=-2.58

Conclusion of the test:

z stat<z crit

Fail to reject H0.

Do not reject H0.the data do not provide sufficient evidence to conclude that p<0.4

ANSWER:

OPTION C